The table below represents the number of republicans, democrats and independent voting on a specific issue. A person from the group is randomly selected. What is the probability of

i. The person being a democrat.  

ii. The person being against the issue.

iii. The person being against the issue OR being a republican.  

iv. The person being not concerned and an independent.  

Republicans Democrats Independents
FOR 55 40 25
AGAINST 25 30 20
NOT Concerned 20 10 5
Answer #1

Solution :

Total = 55 + 40 + 25 + 25 + 30 + 20 + 20 + 10 + 5 = 230

i)

P(The person being a democrat) = (40 + 30 + 10) / 230 = 80 / 230 = 0.3478

ii)

P(The person being against the issue) = (25 + 30 + 20) / 230 = 75 / 230 = 0.3261

iii)

P(The person being against the issue OR being a republican)

= (25 + 30 + 20) + (55 + 25 + 20) - 25 / 230

= 150 / 230

= 0.6521

Iv)

P(The person being not concerned and an independent) = 5 / 230 = 0.0217